OpenStudy (anonymous):

Sirus wrote a check for $67. He subtracted that amount form his account balance and found that the balance was $329 after writing the check. Write and solve an equation to find his balance before writing the check.

OpenStudy (anonymous):

Balance before writing was $x x - 67 = 329 is the equation Add 67 x = 329 +67 = 396 balance before writing the check = $ 396